-62.5=20log(v/0.775)

-62.5=20log(v/0.775)
Answer #1

I think you asked to find v

so we are going to use log property

-62.5=20log(v/0.775)

log(v/0.775)=-62.5/20  

log(v/0.775)=-3.125 {because 62.5/20=3.125}

v/0.775=10^(-3.125) {use the log property log10(x) = y then x=10^(y))

v=0.775*10^(-3.125)

now you can calculate it from calculator by entering power of 10 is -3.125 and multiply with 0.775

than v=5.8116*10^(-4)

I HOPE YOU WILL GET THIS IF YOU WANT SOMETHING ELSE YOU CAN LET ME KNOW IN COMMENTS SECTION